# SESR-Eval: Dataset for Evaluating LLMs in the Title-Abstract Screening of Systematic Reviews

### v1

@@ V-A Creating the benchmark dataset
Some prior studies [66] have bypassed the challenges associated with research artifacts. They reused the original authors’ search queries, re-executed them, and then treated the studies explicitly included in the published review as the included set, with the remainder of the query results inferred as excluded studies. While this approach simplifies the data collection, it does not accurately replicate the original review process. It fails to capture the accurate set of studies that were excluded during the review. It also does not capture the phase in which a paper was excluded, e.g., a paper might be included based on title-abstract screening and later excluded in the full-paper screening phase. Finally, search queries are difficult to replicate retrospectively, as academic databases are continuously updated. Thus, this method introduces uncertainty in identifying excluded studies.
@@ V-B Screening performance - LLMs vs secondary studies
While the seven best performing LLMs we tested offered similar performance, no LLM reached a high recall while maintaining reasonable precision. We suspect this is a limitation of the current LLM screening architecture. The two smallest models (GPT-4.1 nano and Ministral 8B) underperformed relative to the larger LLMs. This suggests that, once a suitable model size is reached, the specific choice of LLM is less critical than the fit between LLM-based screening and the characteristics of the secondary study being screened. A somewhat similar finding has been reported in the biomedical domain in the study by Dennstadt [6]. They screened 38,426 primary studies from 10 secondary studies and observed that performance varied depending on both the LLM and the secondary study used, which aligns with our observations.
Dennstadt [6] also tested the use of a Likert scale in screening, which had a notable impact on performance. We similarly observed that choosing a different Likert scale point for the inclusion boundary had a notable effect and introduced variation between studies.
Work by Delgado [8] represents the largest prior study in terms of the number of LLMs evaluated (18), although they included only three secondary studies. In the medical domain, they reported screening accuracies of 0.920.920.92, 0.880.880.88, and 0.400.400.40 for their three secondary studies. Our results are closely aligned with Delgado’s, as the screening accuracy across our secondary studies ranged from 0.340.340.34 to 0.750.750.75 (Table IX).
The related work in SE [3, 4] reported similar performance for LLMs as we did. However, they only evaluated a single LLM with two studies [4], and two LLMs with a single SR [3]. Thus, our numbers in terms of SRs and LLMs far exceed those. The first work [3] reported a precision of 0.500.500.50 and recall of 0.420.420.42 when screening 1,306 primary studies from a single secondary study using GPT-4. Compared to our results, we surpass both the precision and recall with Claude 3.7 Sonnet (0.560.560.56 and 0.460.460.46), while GPT-4.1 mini offers the best precision in our study (0.600.600.60 and 0.430.430.43 for precision and recall). The second work [4] used one LLM and two primary studies and reported accuracies between 0.630.630.63 and 0.860.860.86. Our study exceeds that range, which is not surprising as we had more LLMs and more secondary studies under evaluation.
@@ V-D Time, token efficiency, and cost of the LLMs
From the tested LLMs and their corresponding timings (Table XI), we see that the fastest LLM per paper was GPT-4.1 nano. However, it was the second-worst performing LLM in the screening tasks. This means that it is not advisable to choose the fastest model for title-abstract screening. On the other hand, looking at the highest-scoring LLMs from Table VIII and their corresponding screening performance in Table VII, Llama 4 Maverick is the highest scoring LLM in four categories: (1) Secondary study accuracy, (2) Primary study accuracy, (3) Primary study F1 score and (4) Secondary study F1 score. Looking at the costs, Llama 4 Maverick cost only \$1.9 per secondary study, which is many times cheaper than most of the tested LLMs. This highlights that it is not optimal to always choose the most expensive LLM for title-abstract screening. Yet, it takes 23 seconds for Llama 4 Maverick to screen one primary study, which was much slower than most of the tested LLMs. However, requests to LLMs can be parallelized, which reduces the total time to screen the primary studies substantially. For a balanced selection, it would make sense to choose an LLM that screens papers efficiently; meaning that it screens papers quickly with an acceptable accuracy; and has reasonable costs.
@@ V-E3 Only zero-shot prompting technique tested
As our dataset has over 30,000 primary studies while targeting to cover as much secondary studies as possible, we only ran prompts that do not include task examples (zero-shot), or additional reasoning instructions (chain-of-thought). Although previous studies have experimented on various different prompting techniques (e.g. one-shot, few-shot, chain-of-thought) [3], we believe that the reasoning LLMs, which in our case were o3-mini and DeepSeek R1, are sufficient to overcome this limitation. Testing different prompting techniques is a feasible area for future research that our dataset enables.
@@ Acknowledgment
The research artifact for the study is available in Zenodo [33].

### v2

@@ V-A Creating the benchmark dataset
Some prior studies [kusa2023csmed] have bypassed the challenges associated with research artifacts. They reused the original authors’ search queries, re-executed them, and then treated the studies explicitly included in the published review as the included set, with the remainder of the query results inferred as excluded studies. While this approach simplifies the data collection, it does not accurately replicate the original review process. It fails to capture the accurate set of studies that were excluded during the review. It also does not capture the phase in which a paper was excluded, e.g., a paper might be included based on title-abstract screening and later excluded in the full-paper screening phase. Finally, search queries are difficult to replicate retrospectively, as academic databases are continuously updated. Thus, this method introduces uncertainty in identifying excluded studies.
@@ V-B Screening performance - LLMs vs secondary studies
While the seven best performing LLMs we tested offered similar performance, no LLM reached a high recall while maintaining reasonable precision. We suspect this is a limitation of the current LLM screening architecture. The two smallest models (GPT-4.1 nano and Ministral 8B) underperformed relative to the larger LLMs. This suggests that, once a suitable model size is reached, the specific choice of LLM is less critical than the fit between LLM-based screening and the characteristics of the secondary study being screened. A somewhat similar finding has been reported in the biomedical domain in the study by Dennstadt [dennstadtTitleAbstractScreening2024]. They screened 38,426 primary studies from 10 secondary studies and observed that performance varied depending on both the LLM and the secondary study used, which aligns with our observations.
Dennstadt [dennstadtTitleAbstractScreening2024] also tested the use of a Likert scale in screening, which had a notable impact on performance. We similarly observed that choosing a different Likert scale point for the inclusion boundary had a notable effect and introduced variation between studies.
Work by Delgado [delgado-chavesTransformingLiteratureScreening2025] represents the largest prior study in terms of the number of LLMs evaluated (18), although they included only three secondary studies. In the medical domain, they reported screening accuracies of $0.92$, $0.88$, and $0.40$ for their three secondary studies. Our results are closely aligned with Delgado’s, as the screening accuracy across our secondary studies ranged from $0.34$ to $0.75$ (Table IX).
The related work in SE [huotalaPromiseChallengesUsing2024a, felizardoChatGPTApplicationSystematic2024] reported similar performance for LLMs as we did. However, they only evaluated a single LLM with two studies [felizardoChatGPTApplicationSystematic2024], and two LLMs with a single SR [huotalaPromiseChallengesUsing2024a]. Thus, our numbers in terms of SRs and LLMs far exceed those. The first work [huotalaPromiseChallengesUsing2024a] reported a precision of $0.50$ and recall of $0.42$ when screening 1,306 primary studies from a single secondary study using GPT-4. Compared to our results, we surpass both the precision and recall with Claude 3.7 Sonnet ($0.56$ and $0.46$), while GPT-4.1 mini offers the best precision in our study ($0.60$ and $0.43$ for precision and recall). The second work [felizardoChatGPTApplicationSystematic2024] used one LLM and two primary studies and reported accuracies between $0.63$ and $0.86$. Our study exceeds that range, which is not surprising as we had more LLMs and more secondary studies under evaluation.
@@ V-D Time, token efficiency, and cost of the LLMs
From the tested LLMs and their corresponding timings (Table XI), we see that the fastest LLM per paper was GPT-4.1 nano. However, it was the second-worst performing LLM in the screening tasks. This means that it is not advisable to choose the fastest model for title-abstract screening. On the other hand, looking at the highest-scoring LLMs from Table VIII and their corresponding screening performance in Table VII, Llama 4 Maverick is the highest scoring LLM in two categories: (1) Secondary study accuracy and (2) Primary study F1 score. Looking at the costs, Llama 4 Maverick cost only \$1.9 per secondary study, which is many times cheaper than most of the tested LLMs. This highlights that it is not optimal to always choose the most expensive LLM for title-abstract screening. Yet, it takes 23 seconds for Llama 4 Maverick to screen one primary study, which was much slower than most of the tested LLMs. However, requests to LLMs can be parallelized, which reduces the total time to screen the primary studies substantially. For a balanced selection, it would make sense to choose an LLM that screens papers efficiently; meaning that it screens papers quickly with an acceptable accuracy; and has reasonable costs.
@@ V-E3 Only zero-shot prompting technique tested
As our dataset has over 30,000 primary studies while targeting to cover as much secondary studies as possible, we only ran prompts that do not include task examples (zero-shot), or additional reasoning instructions (chain-of-thought). Although previous studies have experimented on various different prompting techniques (e.g. one-shot, few-shot, chain-of-thought) [huotalaPromiseChallengesUsing2024a], we believe that the reasoning LLMs, which in our case were o3-mini and DeepSeek R1, are sufficient to overcome this limitation. Testing different prompting techniques is a feasible area for future research that our dataset enables.
@@ Acknowledgment
The research artifact for the study is available in Zenodo [zenodo_2_anon].